# Chaotic-Integrable Transition for Disordered Orbital Hatsugai-Kohmoto Model

### v1

@@
We have drawn connections between the Sachdev-Ye-Kitaev model and the multi-orbit Hatsugei-Kohmoto model, emphasizing their similarities and differences regarding chaotic behaviors. The features of the spectral form factor, such as the dip-ramp-plateau structure and the adjacent gap ratio, indicate chaos in the disordered orbital Hatsugei-Kohmoto model. One significant conclusion is that the plateau value of the out-of-time-order correlator, whether in the Hatsugei-Kohmoto model, Sachdev-Ye-Kitaev model with two- or four-body interactions, or a disorder-free Sachdev-Ye-Kitaev model, does not effectively differentiate between integrable and chaotic phases in many-body systems. This observation suggests a limitation in using out-of-time-order correlator plateau values as a diagnostic tool for chaos. Our exploration of these ideas provides a deeper understanding of how chaos arises in non-Fermi liquid systems and the tools we use to study it. It opens the door to further questions, particularly about whether there are more effective ways to distinguish between chaotic and integrable phases in these complex systems.
@@ 1 Introduction
Non-Fermi liquids (NFLs) are characterized by the absence of well-defined quasiparticles, distinguishing them from Fermi liquids. The notable NFL models such as the Sachdev-Ye-Kitaev (SYK) model [1, 2] and the multi-channel Kondo impurity (MCKI) model, both of which exhibit chaotic behavior [2, 3, 4, 5].
Recently, renewed interest has emerged in a class of exactly solvable many-body fermionic models with infinite-range interactions, initially proposed by Hatsugai and Komohto (HK) [6]. It has been found that the ground states of the HK model do not exhibit the characteristics of Fermi liquids. By introducing an energy cost $U$ for double occupancy in each momentum state, the HK model can be expressed as:
@@
Here $\varepsilon_{k}$ denotes the dispersion in a tight binding model with no orbital degrees of freedom, with the band index $k$ in the one-dimensional Brillouin zone (BZ), and $\sigma=\uparrow,\downarrow$ indicates the $z$-component of the electron spin. The number operator is defined as
$n_{\alpha,\sigma}=c^{\dagger}_{\alpha,\sigma}c_{\alpha,\sigma}$ (2)
under the many-body basis. This paper considers HK-type interactions that impose an energy penalty on doubly-occupied orbital degrees of freedom $\alpha$. An extension of the HK model, which includes an orbital term [7]
$H_{\mathrm{OHK}}=\sum_{k,\alpha<\alpha^{\prime},\sigma}{\big{(}t_{\alpha\alpha^{\prime}\sigma}(k)c^{\dagger}_{k\alpha\sigma}c_{k\alpha^{\prime}\sigma}}+\mathrm{H.C.}\big{)}-\mu\sum_{k,\alpha}{n_{k\alpha\sigma}}+\sum_{k,\alpha,\alpha^{\prime}}{U_{\alpha\alpha^{\prime}}(k)n_{k\alpha\uparrow}n_{k\alpha^{\prime}\downarrow}},$ (3)
where we assume
$U_{\alpha,\alpha^{\prime}}(k)=U_{k,\alpha}\delta_{\alpha,\alpha^{\prime}}$ (4)
respecting the Pauli principle. The chemical potential is set to zero for simplicity, i.e., $\mu=0$. We adjust the permissible orbital number in the half-filling scenario with zero total spin. Extending the HK model to include orbital degrees of freedom preserves its non-fermi liquid behavior. The solvability of the HK model arises from its locality in momentum space. However, when viewed in position space, the interaction term becomes highly non-local, coupling position-space points through a four-point function. This is similar to the SYK model with a four-body interaction (SYK4 model). Exploring the chaotic phenomena of the orbital HK model is significant.
@@
For example, for the Gaussian orthogonal ensemble (GOE), the level spacing follows
$P(s)=\frac{\pi s}{2}\exp\bigg{(}-\frac{\pi s^{2}}{4}\bigg{)},$ (5)
where $s$ is the normalized spacing between neighboring energy levels. This distribution shows a low probability of two energy levels being very close together or far apart, with a peak indicating some typical spacing between levels.
The distribution of energy levels provides the uniquely chaotic properties in a many-body quantum system. Currently, we lose a general method to study quantum chaos in a few-body system [15]. Hence, diagnosing quantum chaos is still challenging. Another diagnostic method from the energy spectrum is the spectral form factor (SFF), which is the square of the Fourier transform of the empirical spectral density [16, 17]. With the disorder average, the SFF ($g$), the disconnected part ($g_{d}$) and the connected part ($g_{c}$) are defined:
@@
In chaotic quantum systems, the SFF is expected to display the characteristic ”dip-ramp-plateau” structure [16, 17, 18]. Initially, the SFF decreases (dip) due to interference between energy levels. As the system evolves, the SFF ramps up, reflecting the emergence of a correlation between energy levels. Eventually, it reaches a plateau, indicating the saturation of correlations, a feature of quantum chaos. The disorder parameter or randomness can generate the shape of the SFF in an integrable system [19, 20]. Therefore, the SFF’s shape alone is insufficient to uniquely diagnose quantum chaos [19, 20]. This limitation suggests that other spectral observables or diagnostics are necessary to capture the few-body systems’ chaotic dynamics fully.
@@
where $W(t)$ and $V(t)$ are operators in the Heisenberg representation, and the angle bracket denotes the thermal average are a diagnostic tool for detecting the chaotic features in quantum systems. These correlation functions reflect how operators evolve and fail to commute over time. In the context of chaos, OTOCs show exponential growth, a hallmark of chaotic behavior. In the semi-classical limit, OTOCs exhibit exponential growth up to a time scale known as the Ehrenfest time. The growth rate is characterized by the Lyapunov exponent $\lambda_{L}$ (the exponent of the OTOCs when it is non-negative), which measures how rapidly initial perturbations grow over time. In some cases, such as the SYK model, the Lyapunov exponent reaches a theoretical upper bound [2, 22]. The disorder-free SYK model [20] though not having random matrix behavior, still shows chaotic characteristics with a non-zero Lyapunov exponent when the system size is large enough [23]. Introducing disorder into systems can induce chaotic behavior. However, it is essential to differentiate this induced randomness from inherent quantum chaos [19, 20]. The presence of randomness does not necessarily mean that the system is chaotic. In general, exchanging the logarithm and the thermal average influences the Lyapunov exponent [24]. Therefore, the Lyapunov exponent generally depends on regularization, unless considering a many-body system. The late-time behavior of OTOCs, precisely the plateau value of the correlation function, is a topic of ongoing investigation [25]. It could serve as an indicator of chaos in general systems, but questions remain about whether this plateau is merely a finite-size artifact or holds deeper physical meaning.
@@
This paper examines chaotic behaviors in non-Fermi liquid systems, specifically focusing on the disordered orbital HK model. By introducing disorder to the model as that $t_{\alpha\alpha^{\prime}\sigma}(k)$ ($t_{h}$) and $U_{k,\alpha}$ ($U$) follow the Gaussian distribution with the zero mean and the variance of $t_{h}$ and $U$ can be tuned and studying statistical distributions, the paper provides insights into how chaos emerges. After the Fourier transformation to the position space, the disordered orbital HK model is similar to the SYK2+SYK4 model. Key points include using the adjacent gap ratio and spectral form factor to characterize chaos and transition between integrable and chaotic phases. The paper also calculates OTOCs to explore the impact of regularization on late-time behavior by comparing the disordered orbital HK model to the SYK2, SYK4, and disordered-free SYK models. Our paper enhances our understanding of chaos in disordered non-Fermi liquids, particularly through the novel application of statistical tools like the adjacent gap ratio, the SFF, and OTOCs to probe the late-time scale. To summarize our results:
@@
  • A phase diagram for the variances of the disorder parameters, $t_{h}$ and $U$, is constructed using the adjacent gap ratio in the disordered orbital HK model. This model has the GOE by turning on the four-point term. The SFF is consistent with the presence of chaos in the disordered orbital HK model.
@@ 2 Disordered Orbital HK Model
The Hamiltonian of the disordered orbital HK model, after Fourier transformation
@@
$+\sum_{x_{1,2,3,4}\alpha}\sum_{k}{U_{k\alpha}e^{ik(x_{1}-x_{2}+x_{3}-x_{4})}c^{\dagger}_{x_{1}\alpha\uparrow}c_{x_{2}\alpha\uparrow}c^{\dagger}_{x_{3}\alpha\downarrow}c_{x_{4}\alpha\downarrow}},$
@@
$H_{\mathrm{SYK2+SYK4}}=\frac{i}{2!}\sum^{N}_{i_{1},i_{2}=1}{\kappa_{i_{1}i_{2}}\chi_{i_{1}}\chi_{i_{2}}}+\frac{1}{4!}\sum^{N}_{i_{1},i_{2},i_{3},i_{4}=1}{J_{i_{1}i_{2}i_{3}i_{4}}\chi_{i_{1}}\chi_{i_{2}}\chi_{i_{3}}\chi_{i_{4}}},$ (14)
where $\chi_{i_{1}}$ represents Majorana fermions, and the random couplings $J_{ijkl}$ and $\kappa_{ij}$ follow the Gaussian distribution with the zero mean and the standard deviations $\sqrt{6}J/N^{3/2}$ and $\kappa/\sqrt{N}$, respectively. We can generally set the fixed-coupling constant $J$ as one because the results only depend on the ratio of two fixed coupling constants $J$ and $\kappa$. We define $N$ as the number of Majorana fermion fields. The two-body interaction term is called the SYK2 model, and another four-body interaction is called the SYK4 model. A chaotic-integrable transition has been observed in the SYK2+SYK4 model [10]. This raises the question of whether a similar transition might occur in the disordered orbital HK model.
Before proceeding with our chaos analysis, we first summarize the known, unknown results, and our results in the non-Fermi liquid models and the related models that we will consider in the tabel 1.
@@ 3.1 Adjacent Gap Ratio
For a quantum chaotic system, the energy levels repel each other, resulting in a probability distribution that follows Wigner-Dyson (WD) statistics [12, 13],
@@
see Figs. 1(a) and 1(b). In the figures, we remove the subscript $n$ for $r_{n}$ for the convenient reading. As the system approaches chaotic behavior, the distribution tends to resemble the GOE distribution, consistent with our Hamiltonian preserving real symmetric properties.
Figure 1: Logarithmic ratio distribution of $\ln r$ over 100 random samples for the disordered orbital HK model with 8 orbitals with half-filling and 0 total spin. (a) Chaotic phase with $\langle\tilde{r}\rangle=0.530$, $\langle t_{h}^{2}\rangle=1$, and $\langle U^{2}\rangle=3.3$. (b) Integrable phase with $\langle\tilde{r}\rangle=0.387$, $\langle t_{h}^{2}\rangle=1$, and $\langle U^{2}\rangle=0$. (c) Phase diagram of the average adjacent gap ratio as a function of $t_{h}$ and $U$ (varied in 0.5 units) over 50 random samples.
@@
$\langle\tilde{r}\rangle\approx 0.387$ (23)
@@
$\langle\tilde{r}\rangle\approx 0.531.$ (24)
As shown in Fig. 1(c), the phase diagram of $\langle\tilde{r}_{n}\rangle$ indicates the Poisson and GOE statistics with different variance of $t_{h}$ and $U$. When the variance of $U$ tends to zero, the system becomes integrable. When $\langle t_{h}^{2}\rangle/\langle U^{2}\rangle\approx 1/3.3$, the system tends towards GOE. These results indicate that the distribution aligns with RMT predictions. Moreover, as seen in Fig. 2, we compute a system’s average adjacent gap ratio by allowing 2, 4, 6, and 8 orbital numbers. We observe that the integrable regime vanishes, and the $\langle\tilde{r}\rangle$ value approaches the GOE value if we turn on the four-point term when the allowable orbital number becomes more considerable. A subtle distinction arises between the vanishing limit of the variance of $U$ and $t_{h}$. When allowing infinite orbital numbers, while the system reaches the integrable regime as the variance of $U$ vanishes first, it enters the GOE regime when the variance of $t_{h}$ approaches zero first.
Figure 2: Phase diagrams showing the average adjacent gap ratio as a function of $t_{h}$ and $U$, both varied in 0.5 unit intervals: (a) 2 orbitals with results averaged over 2000 random samples; (b) 4 orbitals with 1000 random samples; (c) 6 orbitals with 100 random samples; (d) 8 orbitals with 50 random samples.

### v2

@@
We have drawn connections between the Sachdev-Ye-Kitaev model and the multi-orbit Hatsugai-Kohmoto model, emphasizing their similarities and differences regarding chaotic behaviors. The features of the spectral form factor, such as the dip-ramp-plateau structure and the adjacent gap ratio, indicate chaos in the disordered orbital Hatsugai-Kohmoto model. One significant conclusion is that the plateau value of the out-of-time-order correlator, whether in the Hatsugai-Kohmoto model, Sachdev-Ye-Kitaev model with two- or four-body interactions, or a disorder-free Sachdev-Ye-Kitaev model, does not effectively differentiate between integrable and chaotic phases in many-body systems. This observation suggests a limitation in using out-of-time-order correlator plateau values as a diagnostic tool for chaos. Our exploration of these ideas provides a deeper understanding of how chaos arises in non-Fermi liquid systems and the tools we use to study it. It opens the door to further questions, particularly about whether there are more effective ways to distinguish between chaotic and integrable phases in these complex systems.
@@ 1 Introduction
Non-Fermi liquids (NFLs) are characterized by the absence of well-defined quasiparticles, which distinguishes them from Fermi liquids. The notable NFL models, such as the Sachdev-Ye-Kitaev (SYK) model [1, 2] and the multi-channel Kondo impurity (MCKI) model, both of which exhibit chaotic behavior [2, 3, 4, 5]. Recently, renewed interest intertwining of topology and correlations has emerged in a class of exactly solvable many-body fermionic models with infinite-range interactions [6], initially proposed by Hatsugai and Komohto (HK) [7]. These models are described by topologically nontrivial band topology together with the HK interactions, where topological Mott insulating phases emerge [8, 6]. It has been found that the ground states of the HK model do not exhibit the characteristics of Fermi liquids. By introducing an energy cost $U$ for double occupancy in each momentum state, the HK model can be expressed as
@@
Here $\varepsilon_{k}$ denotes the dispersion in a tight binding model with no orbital degrees of freedom, with the band index $k$ in the one-dimensional Brillouin zone (BZ), and $\sigma=\uparrow,\downarrow$ indicates the $z$-component of the electron spin. The number operator is defined as $n_{\alpha,\sigma}=c^{\dagger}_{\alpha,\sigma}c_{\alpha,\sigma}$ under the many-body basis. This paper considers HK-type interactions that impose an energy penalty on doubly-occupied orbital degrees of freedom $\alpha$. An extension of the HK model, which includes an orbital term [6]
$H_{\mathrm{OHK}}=$ $\sum_{k,\alpha<\alpha^{\prime},\sigma}{\big{(}t_{\alpha\alpha^{\prime}\sigma}(k)c^{\dagger}_{k\alpha\sigma}c_{k\alpha^{\prime}\sigma}}+\mathrm{H.C.}\big{)}$
$-\mu\sum_{k,\alpha}{n_{k\alpha\sigma}}+\sum_{k,\alpha,\alpha^{\prime}}{U_{\alpha\alpha^{\prime}}(k)n_{k\alpha\uparrow}n_{k\alpha^{\prime}\downarrow}},$ (2)
where we assume $U_{\alpha,\alpha^{\prime}}(k)=U_{k,\alpha}\delta_{\alpha,\alpha^{\prime}}$ respecting the Pauli principle. The chemical potential is set to zero for simplicity, i.e., $\mu=0$. We adjust the permissible orbital number in the half-filling scenario with zero total spin. Extending the HK model to include orbital degrees of freedom preserves its non-Fermi liquid behavior. The solvability of the HK model arises from its locality in momentum space. However, when viewed in position space, the interaction term becomes highly non-local, coupling position-space points through a four-point function. This is similar to the SYK model with a four-body interaction (SYK4 model). Exploring the chaotic phenomena of the orbital HK model is significant.
@@
For example, for the Gaussian orthogonal ensemble (GOE), the level spacing follows $P(s)=\frac{\pi s}{2}\exp\bigg{(}-\frac{\pi s^{2}}{4}\bigg{)}$, where $s$ is the normalized spacing between neighboring energy levels. This distribution shows a low probability of two energy levels being very close together or far apart, with a peak indicating some typical spacing between levels.
The distribution of energy levels provides the uniquely chaotic properties in a many-body quantum system. Currently, we lack a general method to study quantum chaos in a few-body system [16]. Hence, diagnosing quantum chaos is still challenging. Another diagnostic method from the energy spectrum is the spectral form factor (SFF), which is the square of the Fourier transform of the empirical spectral density [17, 18]. With the disorder average, the SFF ($g$), the disconnected part ($g_{d}$) and the connected part ($g_{c}$) are defined:
@@
In chaotic quantum systems, the SFF is expected to display the characteristic ”dip-ramp-plateau” structure [17, 18, 19]. Initially, the SFF decreases (dip) due to interference between energy levels. As the system evolves, the SFF ramps up, reflecting the emergence of a correlation between energy levels. Eventually, it reaches a plateau, indicating the saturation of correlations, a feature of quantum chaos. The disorder parameter or randomness can generate the dip-ramp-plateau to the shape of the SFF in an integrable system [20, 21]. Therefore, the SFF’s shape alone cannot uniquely diagnose quantum chaos [20, 21]. This limitation suggests that other spectral observables or diagnostics are necessary to capture the chaotic dynamics of few-body systems fully.
@@
where $W(t)$ and $V(t)$ are operators in the Heisenberg representation, and the angle bracket denotes the thermal average, a diagnostic tool for detecting the chaotic features in quantum systems. These correlation functions reflect how operators evolve and fail to commute over time. In the context of chaos, OTOCs show exponential growth, a hallmark of chaotic behavior. In the semi-classical limit, OTOCs exhibit exponential growth up to a time scale known as the Ehrenfest time. The growth rate is characterized by the Lyapunov exponent $\lambda_{L}$ (the exponent of the OTOCs when it is non-negative), which measures how rapidly initial perturbations grow over time. In some cases, such as the SYK model, the Lyapunov exponent reaches a theoretical upper bound [2, 23]. The disorder-free SYK model [21], though, does not have random matrix behavior, still shows chaotic characteristics with a non-zero Lyapunov exponent when the system size is large enough [24]. Introducing disorder into systems can induce chaotic behavior. However, it is essential to differentiate this induced randomness from inherent quantum chaos [20, 21]. The presence of randomness does not necessarily mean that the system is chaotic. In general, exchanging the logarithm and the thermal average influences the Lyapunov exponent [25]. Therefore, the Lyapunov exponent generally depends on regularization, unless considering a many-body system. The late-time behavior of OTOCs, precisely the plateau value of the correlation function, is a topic of ongoing investigation [26]. It could serve as an indicator of chaos in general systems. However, questions remain about whether this plateau is merely a finite-size artifact or holds deeper physical meaning.
@@
This paper examines chaotic behaviors in non-Fermi liquid systems, specifically focusing on the disordered orbital HK model. By introducing disorder to the model as that $t_{\alpha\alpha^{\prime}\sigma}(k)$ ($t_{h}$) and $U_{k,\alpha}$ ($U$) follow the Gaussian distribution with the zero mean and the variance of $t_{h}$ and $U$ can be tuned and studying statistical distributions, the paper provides insights into how chaos emerges. After the Fourier transformation to the position space, the disordered orbital HK model is similar to the SYK2+SYK4 model. If the range of the band index goes to infinity, the summation of the momentum index vanishes in the HK model [27]. Hence, this model is simpler. Key points include using the adjacent gap ratio and spectral form factor to characterize chaos and transition between integrable and chaotic phases. The paper also calculates OTOCs to explore the impact of regularization on late-time behavior by comparing the disordered orbital HK model to the SYK2, SYK4, and disordered-free SYK models. Our paper elucidates quantum chaotic behaviors in disordered non-Fermi liquids, particularly through the novel application of statistical tools like the adjacent gap ratio, the SFF, and OTOCs to probe the late-time scale. We summarize our results as follows:
@@
  • A phase diagram for the variances of the disorder parameters, $t_{h}$ and $U$, is constructed using the adjacent gap ratio in the disordered orbital HK model. This model has the GOE by turning on the four-point term. The SFF is consistent with chaos in the disordered orbital HK model.
@@ 2 Disordered Orbital HK Model
The Hamiltonian of the disordered orbital HK model, after the Fourier transformation
@@
$+\sum_{x_{1,2,3,4}\alpha}\sum_{k}U_{k\alpha}e^{ik(x_{1}-x_{2}+x_{3}-x_{4})}$
×$\times c^{\dagger}_{x_{1}\alpha\uparrow}c_{x_{2}\alpha\uparrow}c^{\dagger}_{x_{3}\alpha\downarrow}c_{x_{4}\alpha\downarrow},$
@@
$H_{\mathrm{SYK2+SYK4}}=$ $\frac{i}{2!}\sum^{N}_{i_{1},i_{2}=1}{\kappa_{i_{1}i_{2}}\chi_{i_{1}}\chi_{i_{2}}}$
$+\frac{1}{4!}\sum^{N}_{i_{1},i_{2},i_{3},i_{4}=1}{J_{i_{1}i_{2}i_{3}i_{4}}\chi_{i_{1}}\chi_{i_{2}}\chi_{i_{3}}\chi_{i_{4}}},$ (11)
where $\chi_{i_{1}}$ represents Majorana fermions, and the random couplings $J_{ijkl}$ and $\kappa_{ij}$ follow the Gaussian distribution with the zero mean and the standard deviations $\sqrt{6}J/N^{3/2}$ and $\kappa/\sqrt{N}$, respectively. We can generally set the fixed-coupling constant $J$ as one because the results only depend on the ratio of two fixed coupling constants $J$ and $\kappa$. We define $N$ as the number of Majorana fermion fields. The two-body interaction term is called the SYK2 model, and another four-body interaction is called the SYK4 model. A chaotic-integrable transition has been observed in the SYK2+SYK4 model [11]. This raises the question of whether a similar transition might occur in the disordered orbital HK model. Comparing the SYK2 + SYK4 [Eq. (11)] and the orbital HK model [Eq. (9)], the former conserves fermionic parity, while the latter conserves charge. We expect that a charge-conserving variant of the SYK model, in which Majorana fermions are replaced by complex fermions, should yield qualitatively similar results to the orbital HK model.
$J_{i_{1}i_{2}i_{3}i_{4}}=1.$ (12)
Before proceeding with our analysis, we first summarize the known and unknown results, and our results in the non-Fermi liquid models and the related models in Tab. 1. The ”$\beta$ ind.” means that the behavior of late-time OTOCs is temperature independent. For the SYK2 and SYK4 models, the analyses are consistent across studies [4, 11, 21]. However, there is a mismatch in the disorder-free SYK2 model between spectral statistics and OTOC analysis [24]. This indicated that even in the absence of random matrix theory properties, the system can still exhibit exponential growth in the early-time behavior of OTOCs [24]. Besides, for another non-Fermi liquid model, the multi-channel Kodon model, exponential growth in early-time OTOCs was observed in the large-$N$ limit [5], but its spectral statistics remain unknown. For our disordered orbital HK model, chaotic behavior is evident from the adjacent gap ratio and SFF. Lastly, since the late-time OTOCs of SYK2, SYK4, and disorder-free SYK2 are temperature-independent, we argue that the temperature dependence observed in late-time behavior arises from the finite size effect and is not a reliable indicator for many-body chaos [26].
@@ 3.1 Adjacent Gap Ratio
For a quantum chaotic system, the energy levels repel each other, resulting in a probability distribution that follows Wigner-Dyson (WD) statistics for $3\times 3$ matrices [13, 14],
@@
see Fig. 1. In the figure, we remove the subscript $n$ for $r_{n}$ for the convenience of reading. As the system approaches chaotic behavior, the distribution tends to resemble the GOE distribution, consistent with our Hamiltonian preserving real symmetric properties.
Figure 1: Logarithmic ratio distribution of $\ln r$ over 5 random samples for the disordered orbital HK model with 10 orbitals with half-filling and 0 total spin. (a) Chaotic phase with $\langle\tilde{r}\rangle=0.531$, $\langle t_{h}^{2}\rangle=1$, and $\langle U^{2}\rangle=3.3$. (b) Integrable phase with $\langle\tilde{r}\rangle=0.386$, $\langle t_{h}^{2}\rangle=1$, and $\langle U^{2}\rangle=0$. (c) The phase diagram of the average adjacent gap ratio as a variance of $U$ over a variance of $t_{h}$. 6 orbitals with 5000 random samples; 8 orbitals with 100 random samples; 10 orbitals with 5 random samples.
@@
$\langle\tilde{r}\rangle\approx 0.386$ (20)
@@
$\langle\tilde{r}\rangle\approx 0.536.$ (21)
As shown in Fig. 1, the phase diagram of $\langle\tilde{r}_{n}\rangle$ indicates the transition between Poisson and GOE statistics for different values of the variance ratio $\langle U^{2}\rangle/\langle t_{h}^{2}\rangle$ in the 6-, 8-, and 10-orbital systems. When this ratio approaches zero, the system becomes integrable. However, for $\langle U^{2}\rangle/\langle t_{h}^{2}\rangle>0.33$ in the 10-orbital case, the system exhibits GOE statistics, aligns with RMT predictions. Moreover, we observe that the number of orbitals increases, the integrable regime tends to vanish, and the value of $\langle\tilde{r}\rangle$ approaches the GOE limit upon introducing the four-point interaction term. In the thermodynamic limit, the system initially enters the integrable regime when $\langle U^{2}\rangle$ vanishes first, but transitions into the chaotic regime if $\langle t_{h}^{2}\rangle$ approaches zero first. Thus, our key conclusion is that a non-zero variance of $U$ drives the system toward chaotic behavior. This finding highlights the crucial role of interaction disorder in inducing quantum chaos within the orbital HK model.